# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

### v1

@@ Abstract
Fully decentralized training of machine learning models offers significant advantages in scalability, robustness, and fault tolerance. However, achieving differential privacy (DP) in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes. In this work, we present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, both with and without secure summation over each node’s direct neighbors. Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage across these scenarios. This framework significantly improves upon prior analyses, for example, reducing the Rényi DP parameter growth from $O(T^{2})$ to $O(T)$, where $T$ is the number of training rounds. We validate our analysis with numerical results demonstrating superior DP bounds compared to existing approaches. We further illustrate our analysis with a logistic regression experiment on MNIST image classification in a fully decentralized setting, demonstrating utility comparable to central aggregation methods.
@@ 1 Introduction
In random walk-based methods (Lopes and Sayed,, 2007; Johansson et al.,, 2009; Mao et al.,, 2020), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix. The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor. Over time, and under mild assumptions on the transition matrix and the network graph (e.g., connectedness, symmetry), the random walk ensures uniform coverage of all nodes. However, such sequential update dynamics inherently limits parallelism and scalability.
@@
In gossip averaging-based methods (Boyd et al.,, 2006; Dimakis et al.,, 2010), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states. Each node updates its parameters by averaging over the received messages. These operations are repeated iteratively until the network reaches approximate consensus on the model parameters. Gossip protocols are naturally parallelizable and more scalable than random walk approaches, making them attractive for large-scale decentralized learning. This work focuses on such gossip-based protocols for DP decentralized learning.
@@
DP in Centralized FL. In centralized FL (Kairouz et al., 2021b, ), distributed DP (Ullah et al.,, 2023; Hartmann and Kairouz,, 2023; Kairouz et al., 2021a, ) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (Truex et al.,, 2019; Erlingsson et al.,, 2019; Bell et al.,, 2020). This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
@@
DP in Decentralized Optimization. The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (Dekker et al.,, 2025; Zhu et al.,, 2019; Mrini et al.,, 2024; Pasquini et al.,, 2023). Early decentralized DP methods achieve privacy by locally perturbing gradients or models (Huang et al.,, 2015; Bellet et al.,, 2018; Xu et al.,, 2022), relying on local DP guarantees. This often imposes a poor utility-privacy tradeoff compared to centralized DP (Chan et al.,, 2012).
@@
To address this, recent works propose pairwise network DP (Cyffers et al.,, 2024, 2022; Cyffers and Bellet,, 2022), where privacy loss is analyzed between pairs of nodes. These relaxations yield better utility while accounting for the structure of decentralized communication. However, for the practical case of gossip averaging, there does not yet exist a satisfactory composition analysis for the pairwise network DP guarantees. To the best of our knowledge, the best bounds in this case are the Rényi differential privacy (RDP) bounds by Cyffers et al., (2022) which exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of training iterations, making them unsuitable for practical ML model training with reasonable privacy guarantees.
@@
Our Approach. In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows empirically as $O(\sqrt{T})$, resulting in $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses. This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs). This also provides tools to analyze the network DP guarantees in different threat scenarios. In addition to considering the threat model of (Cyffers et al.,, 2022), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
@@
Secure Aggregation in Decentralized Learning. While the amplification effect of summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored. In current gossip-based systems (Cyffers et al.,, 2022), each node receives messages directly from its neighbors and can, in principle, view their full content. As we show, summation over a node’s neighborhood can yield a meaningful privacy amplification. Several studies have proposed integrating MPC primitives into decentralized learning workflows (Jayaraman et al.,, 2018; Lian et al.,, 2018; Jeon et al.,, 2021). Concrete protocols for decentralized secure aggregation have been developed in recent works (Sabater et al.,, 2022; Pereira et al.,, 2024; Biswas et al.,, 2024). The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (Nicolas et al.,, 2024).
@@
  • We develop a novel analytical framework for evaluating differential privacy guarantees in decentralized gossip-based averaging algorithms. Our framework leverages a linear dynamical systems perspective to track sensitivity propagation through iterative updates.
  • We demonstrate that the DP guarantees, for both non-adaptive and adaptive compositions-with or without secure summation-can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
  • Our empirical results show that the squared sensitivity scales as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on Rényi differential privacy (RDP).
  • We validate our findings on a decentralized logistic regression task using the MNIST dataset, showing privacy-utility trade-offs comparable to those of central FL.
@@ 2.1 Differential Privacy
With the hockey-stick divergence we can equivalently measure the $(\varepsilon,\delta)$-distance of distributions. For probability distributions $P$ and $Q$, and for $\varepsilon\in\mathbb{R}$, it is defined as $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(P||Q)=\int[P(t)-{\rm e}\hskip 1.0pt^{\varepsilon}Q(t)]_{+}$, where $[z]_{+}=\max\{0,z\}$. The $(\varepsilon,\delta)$-DP guarantees can be then given as follows.
Lemma 2 (Balle et al., 2018).
@@
The results of (Balle and Wang,, 2018) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism. Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by Dong et al., (2022).
Lemma 4 (Dong et al., 2022).
@@ 2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (Golub and Van Loan,, 2013).
@@
2.3 Decentralized Learning and Network DP
We consider a connected graph $G=(V,E)$, with $V$ and $E$ as the nodes and edges of graph, respectively. The graph structure is encoded into an adjacency matrix $A\in\{1,0\}^{n\times n}$, with $A_{ij}=1$ if nodes $i$ and $j$ are connected by an edge and 0 otherwise. We denote the nodes via indices, i.e., the $n$ nodes of the graph are denoted via $[n]=\{1,\ldots,n\}$. The time is divided into discrete intervals. In each round $t\in[T]$, $T\in\mathbb{Z}_{+}$, each node $i$ can exchange a message with its neighboring nodes $N_{i}=\{j|(i,j)\in E\}$. We denote the closed neighborhood of node $i$ by $\bar{N}_{i}=N_{i}\cup\{i\}$.
The total dataset $D$ is partitioned into $n$ disjoint shards with each node $i$ holding a local dataset $D_{i}$. In decentralized learning the goal is to learn a model $\theta$ that minimizes the loss function $f(\theta)=\sum\nolimits_{i=1}^{n}f_{i}(\theta)$, where $f_{i}(\theta)$ denotes the empirical loss of node $i$, $i\in[n]$, and $\theta\in\mathbb{R}^{d}$ denotes the model parameters. At each round $t\in[T]$, the global state variable is denoted $\theta_{t}\in\mathbb{R}^{n}$ with node-wise variables $[\theta_{t}]_{i}$, $i\in[n]$. In our analysis, we consider univariate node-wise variables. Our results are generalizable to higher dimensions via tensorization using Kronecker products.
DP requires defining a neighborhood relation for datasets. In our experiments we adopt a record-level relation which describes the protection for an individual data element in a dataset $D_{i}$ of a given node $i\in[n]$, however we note that our results are applicable to any neighborhood relation. Formally, $D=\cup_{i\in[n]}D_{i}$ and $D=\cup_{i\in[n]}D_{i}$ are adjacent datasets, denoted $D\sim D^{\prime}$, if there exists $i\in[n]$ such that only $D_{i}$ and $D_{i}^{\prime}$ differ. We use $D\sim_{i}D^{\prime}$ to denote that $D$ and $D^{\prime}$ differ only in the data of node $i$.
We define a decentralized mechanism $\mathcal{M}$ as a randomized function that takes as input the total dataset $D=\cup_{i\in[n]}D_{i}$ and outputs all the shared messages. What is seen by whom, is represented by the view. We denote by $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})$ the part of $\mathcal{M}(D)$ visible to the set of nodes $\mathcal{A}$. For a set of nodes $\mathcal{A}=\{i_{1},\ldots,i_{\left|\mathcal{A}\right|}\}$ we define the selector matrix $S(\mathcal{A})\in\mathbb{R}^{\left|\mathcal{A}\right|\times n}$ as $S(\mathcal{A})=\sum_{k=1}^{\left|\mathcal{A}\right|}e_{k}e_{i_{k}}$, where $e_{k}$ is the $k$th standard basis vector of $\mathbb{R}^{\left|\mathcal{A}\right|}$ and $e_{i_{k}}$ the $i_{k}$th standard basis vector of $\mathbb{R}^{n}$.
We say that the node $j$ is $(\varepsilon,\delta)$-DP from point of view of a subset of nodes $\mathcal{A}\subset[n]$, if $D\simeq_{j}D^{\prime}$ and $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\simeq_{(\varepsilon,\delta)}\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$ We also assume that nodes are honest but curious: they adhere to the protocol, but may attempt to infer additional information from the messages they observe. Similarly to (Cyffers et al.,, 2022), we first assume that the noise injected by the observing nodes contributes to the DP guarantees. In Section 3.5, we show how to remove this assumption.
[…]
@@ 3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., Antoulas,, 2005). In particular, we consider the systems without the feedthrough term in which case their dynamics is are described by the equations
@@
We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{N}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[N]$ (see, e.g., Ch. 4 Antoulas,, 2005). In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
This state-space dynamics perspective can be used to analyze different threat models for both non-adaptive and adaptive compositions. To summarize our results, when measuring the DP guarantees for the data of node $j$, the coefficient matrices $(A,B,C)$ are given as follows.
  1. 1.
    Non-adaptive analysis with secure summation from the view of node $i$:
    $(A,B,C)=(W,W,e^{\top}_{i}),$
    where $e_{i}$ denotes the $i$th standard basis (one-hot) vector in $\mathbb{R}^{n}$.
  2. 2.
    Non-adaptive analysis with secure summation and a set of colluding nodes $\mathcal{C}=\{i_{1},\ldots,i_{\left|\mathcal{C}\right|}\}$:
    $(A,B,C)=\big{(}W,W,S(\mathcal{C})\big{)}$
    with the selector matrix $S(\mathcal{C})$.
  3. 3.
    Non-adaptive and adaptive analysis without secure summation:
    $(A,B,C)=(W,I_{n},e_{j}).$
  4. 4.
    Adaptive analysis with secure summation:
    $(A,B,C)=\big{(}W,W,S(\bar{N}_{j})\big{)}$
    with the selector matrix $S(\bar{N}_{j})$.
We next derive these matrices and the corresponding DP guarantees one by one.
3.2 Analysis of Non-Adaptive Compositions with Secure Summation
@@
where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. From the recursion (3.1) it follows that from the point-of-view of node $i$, the view is given by $\mathrm{View}_{\mathcal{M}(D)}(\{i\})=\begin{bmatrix}y_{1}&\ldots&y_{n}\end{bmatrix}^{\top}$, where
$y_{1}$ $=e_{i}^{\top}(Wx_{0}+Wu_{0})$
$y_{2}$ $=e_{i}^{\top}(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1})$
@@
$y_{T}$ $=e_{i}^{\top}(W^{T}x_{0}+\cdots+Wx_{T-1}+W^{T}u_{0}+\cdots+Wu_{T-1}).$
From this representation we get the following.
@@
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $\left|x_{t}^{j}-{x^{\prime}}_{t}^{j}\right|\leq 1$ (this without loss of generality). Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
@@
$\widetilde{x}_{T}=\begin{bmatrix}e_{i}^{\top}We_{j}\\ e_{i}^{\top}(W^{2}+W)e_{j}\\ \vdots\\ e_{i}^{\top}(W^{T-1}+\cdots+W)e_{j}\end{bmatrix}\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}e_{i}^{\top}W&0&0&\dots&0\\ e_{i}^{\top}W^{2}&e_{i}^{\top}W&0&\dots&0\\ e_{i}^{\top}W^{3}&e_{i}^{\top}W^{2}&e_{i}^{\top}W&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ e_{i}^{\top}W^{T}&e_{i}^{\top}W^{T-1}&\dots&e_{i}^{\top}W^{2}&e_{i}^{\top}W\end{bmatrix}$ (3.2)

### v2

@@ Abstract
Fully decentralized training of machine learning models offers significant advantages in scalability, robustness, and fault tolerance. However, achieving differential privacy (DP) guarantees in such settings is challenging due to the absence of a central aggregator and varying trust assumptions among nodes. We present a novel privacy analysis of decentralized gossip-based averaging algorithms with additive node-level noise, from arbitrary views of nodes in a graph and especially consider the averaging over nearest neighbors with secure summation and individual node-wise views. Our main contribution is a new analytical framework based on a linear systems formulation that accurately characterizes privacy leakage between nodes across different scenarios. In case the gossip averaging happens via secure summation, we show that the Rényi DP parameter growth is asymptotically $O(T)$, where $T$ is the number of training rounds, similarly as in the case of central aggregation.
@@ 1 Introduction
In random walk-based methods (rw1; rw2; rw3), a node computes local gradients and sends its model state to a randomly selected neighbor, sampled according to a doubly stochastic mixing matrix. The neighbor then updates its local parameters and passes the updated model to another randomly chosen neighbor. Over time, and under mild assumptions on the transition matrix and the network graph (e.g., connectedness, symmetry), the random walk ensures uniform coverage of all nodes. However, such sequential update dynamics inherently limits parallelism and scalability.
@@
In gossip averaging-based methods (gossip2; gossip1), all nodes simultaneously communicate with their neighbors in synchronous rounds, sharing either gradients or full model states. Each node updates its parameters by averaging over the received messages. These operations are repeated iteratively until the network reaches approximate consensus on the model parameters. Gossip protocols are naturally parallelizable and more scalable than random walk approaches, making them attractive for large-scale decentralized learning. This work focuses on such gossip-based protocols for DP decentralized learning.
@@
DP in Centralized FL. In centralized FL (FLbook), distributed DP (UllahCKO23; google_flblog_2023; FLDP3) strengthens statistical privacy guarantees by combining local noise injection with cryptographic tools such as secure summation (truex2019hybrid; Erlingsson2019; bell2020secure). This not only eliminates the need to trust individual nodes but also enables lower per-user noise by aggregating over multiple contributions.
@@
DP in Decentralized Optimization. The increased communication and lack of central control in decentralized optimization expands the attack surface, exposing the system to more potential privacy breaches (dlattack4; dlattack3; dlattack1; dlattack2). Early decentralized DP methods achieve privacy by locally perturbing gradients or models (HMV2015; BGTT18; XZW2022), relying on local DP guarantees. This often imposes a poor utility-privacy tradeoff compared to centralized DP (ChanSS12).
@@
To address this, recent works propose pair-wise network DP (networkdp3; networkdp2; networkdp1), where privacy loss is analyzed between pairs of nodes. These relaxations give better utility while accounting for the structure of decentralized communication. networkdp2 give Rényi differential privacy (RDP) bounds that take into account the limited view, however they exhibit a $T^{2}$-growth of the RDP parameters, where $T$ is the number of compositions (or training iterations in ML model training), making them unsuitable for practical ML model training with reasonable privacy guarantees. Recently, bellet2025unified considered the linear-systems viewpoint of decentralized optimization to cover adaptive compositions through the so called matrix mechanisms (kairouz2021practical; pillutla2025correlated) which leads to considerably more accurate privacy analysis than that of networkdp2. Nevertheless, as our asymptotic sensitivity result (our Theorem 13) shows, the full linear-systems framework is still a valuable tool for understanding the privacy-preserving properties of decentralized optimization.
@@
Our Approach. In this work, we develop techniques to analyze network DP guarantees under various threat models, showing that the total sensitivity grows asymptotically as $O(\sqrt{T})$, resulting in asymptotic $O(T)$ growth of the RDP parameters, for example, representing a significant improvement over prior analyses. This network DP accounting is obtained by interpreting the dynamics of the gossip averaging as a linear state-space system and the view of individual nodes or subsets of nodes as projected Gaussian mechanisms (GMs). This also provides tools to analyze the network DP guarantees in different threat scenarios. In addition to considering the threat model of (networkdp2), where the neighboring nodes see the plain messages sent by their neighbors, we also analyze algorithms that incorporate a secure summation protocol between neighboring nodes.
@@
Secure Aggregation in Decentralized Learning. While the amplification effect of summation is well-understood in centralized FL, its impact in fully decentralized settings remains underexplored. In current gossip-based systems (networkdp2), each node receives messages directly from its neighbors and can, in principle, view their full content. As we show, summation over a node’s neighborhood can give a meaningful privacy amplification. Several studies have proposed integrating MPC primitives into decentralized learning workflows (jayaraman2018distributed; lian2018asynchronous; jeon2021privacy). Concrete protocols for decentralized secure aggregation have been developed in recent works (sabater2022gopa; dlsa1; biswas2024). The use of secure summation for private decentralized algorithms has also been considered in specific applications such as differentially private PCA (nicolas2024differentially).
@@
  • We provide an analytical framework for evaluating differential privacy guarantees in decentralized gossip-based averaging algorithms. Our framework leverages a linear dynamical systems perspective to track sensitivity propagation through iterative updates.
  • We demonstrate that the DP guarantees can be directly characterized using the Gaussian mechanism, with sensitivity and noise scale computed with the help of our linear system model.
  • We show that in case secure summation is used to carry averaging over the nearest neighbors, the squared sensitivity scales asymptotically as $T$ over $T$ training rounds, significantly improving over the $O(T^{2})$ growth in state-of-the-art analyses based on RDP (networkdp2).
@@ 2.1 Differential Privacy
With the hockey-stick divergence we can equivalently measure the $(\varepsilon,\delta)$-distance of distributions. For probability distributions $P$ and $Q$, and for $\varepsilon\in\mathbb{R}$, it is defined as $H_{{\rm e}\hskip 1.0pt^{\varepsilon}}(P||Q)=\int[P(t)-{\rm e}\hskip 1.0pt^{\varepsilon}Q(t)]_{+}$, where $z]_{+}=\max\{0,z\}$. The $(\varepsilon,\delta)$-DP guarantees can be then given as follows.
Lemma 2 (balle2018subsampling).
@@
The results of (balle2018gauss) give the tight $(\varepsilon,\delta)$-guarantees for the Gaussian mechanism. Bounds for compositions follow from the fact that the Gaussian mechanism is $\mu$-Gaussian Differentially Private for $\mu=\Delta_{2}/\sigma$, and from the composition results for $\mu$-GDP mechanisms by dong2022gaussian.
Lemma 4 (dong2022gaussian).
@@ 2.2 Projected Gaussian Mechanism and Moore–Penrose Pseudoinverse
We consider the computationally tractable definition of the Moore–Penrose pseudoinverse based on the singular value decomposition (SVD) (golub2013matrix).
@@
2.3 Decentralized Learning and Network Differential Privacy
We consider a connected undirected graph $G=(V,E)$ with $|V|=n$ nodes. The graph structure is encoded by an adjacency matrix
$A\in\{0,1\}^{n\times n},\qquad A_{ij}=1\;\text{iff}\;(i,j)\in E.$
We denote the nodes by indices $n]=\{1,\dots,n\}$. For each node $i\in[n]$, the neighborhood and closed neighborhood are defined as
$N_{i}=\{j\in[n]:(i,j)\in E\},\qquad\bar{N}_{i}=N_{i}\cup\{i\}.$
Local Data and Learning Objective.
The global dataset is partitioned across nodes:
$D=\bigcup_{i=1}^{n}D_{i},$
where node $i$ holds the shard $D_{i}$. The goal of decentralized learning is to minimize the global loss
$f(\theta)=\sum_{i=1}^{n}f_{i}(\theta),\qquad\theta\in\mathbb{R}^{d},$
where $f_{i}(\theta)$ depends only on $D_{i}$. At each round $t\in[T]$, node $i$ maintains a state variable $\theta_{t}]_{i}$. For the analysis, we consider the scalar case ($d=1$); the multidimensional case follows from the matrix mechanisms analogy given by bellet2025unified as described in Section 3.4.
Dataset Adjacency.
We adopt a record-level adjacency relation. Two datasets $D=\cup_{i}D_{i}$ and $D^{\prime}=\cup_{i}D_{i}^{\prime}$ are adjacent, written $D\sim D^{\prime}$, if they differ in exactly one node’s data:
$D\sim_{i}D^{\prime}\quad\iff\quad D_{j}=D^{\prime}_{j}\ \text{for all }j\neq i.$
Mechanism and Observations.
A decentralized mechanism is a randomized mapping
$\mathcal{M}:D\mapsto\text{(all messages exchanged during the protocol)}.$
For a subset of nodes $\mathcal{A}\subseteq[n]$, the view of $\mathcal{A}$ is denoted
$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A}).$
To extract node-wise components, we define the selector matrix: for $\mathcal{A}=\{i_{1},\dots,i_{|\mathcal{A}|}\}$,
$S(\mathcal{A})=\sum_{k=1}^{|\mathcal{A}|}e_{k}e_{i_{k}}^{\top}\in\mathbb{R}^{|\mathcal{A}|\times n},$
where $e_{k}$ is the $k$th standard basis vector.
Network Differential Privacy.
Node $j$ is $(\varepsilon,\delta)$-DP from the point of view of $\mathcal{A}\subseteq[n]$ if for all adjacent datasets $D\sim_{j}D^{\prime}$,
$\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})\;\simeq_{(\varepsilon,\delta)}\;\mathrm{View}_{\mathcal{M}(D^{\prime})}(\mathcal{A}).$
We assume honest-but-curious nodes: they follow the protocol but analyze observed messages. As in (networkdp2), we first allow noise added by observers to contribute to privacy; this assumption is removed in Section 3.3.
Gossip-Based Averaging.
Let $x_{t}\in\mathbb{R}^{n}$ denote the vector of local values exchanged at time $t$. A gossip update is given by
$\theta_{t+1}=Wx_{t},$
where $W\in\mathbb{R}^{n\times n}$ is the gossip matrix. We call $W$:
$\text{row-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1},\qquad\text{doubly-stochastic}\;\iff\;W\mathbf{1}=\mathbf{1}\ \text{and}\ W^{\top}\mathbf{1}=\mathbf{1}.$
For the error analysis we also consider the definition of primitive gossip matrices.
Definition 2.1 (Primitive Matrix).
A matrix $W\in\mathbb{R}^{n\times n}$ is called primitive if there exists $k\in\mathbb{N}$ such that
$W^{k}>0\quad\text{(entrywise)}.$
Equivalently, a row-stochastic matrix is primitive if and only if the induced Markov chain is irreducible and aperiodic. In this case, $W$ has a unique stationary distribution $\pi>0$ with $\pi^{\top}W=\pi^{\top}$, and moreover
$W^{t}\to\mathbf{1}\pi^{\top}\qquad\text{as }t\to\infty.$
@@ 3.1 Gossip Averaging as Discrete-Time Linear State-Space Dynamics
We analyze gossip averaging algorithms by viewing their dynamics as those of discrete-time linear state-space systems (see, e.g., antoulas2005approximation). In particular, we consider the systems without the feedthrough term in which case their dynamics is are described by the equations
@@
We utilize techniques of representing the sequence of observations $y_{1},\ldots,y_{N}$ as a large linear system, where $u_{t}$’s are vectorized and the global dynamics is described by a large block-lower-triangular state transition matrix determined by $A_{t}$, $B_{t}$ and $C_{t}$, $t\in[N]$ (see, e.g., Ch. 4 antoulas2005approximation). In our presentation, we focus on time-invariant graphs in which case $A_{t},B_{t}$ and $C_{t}$ are fixed.
3.2 Analysis of Non-Adaptive Compositions
@@
where $u_{t}\sim\mathcal{N}(0,\sigma^{2}I_{n})$. From the recursion (3.1) it follows that from the view determined by the selector $S$, the view is given by $\mathrm{View}_{\mathcal{M}(D)}(\mathcal{A})=\begin{bmatrix}y_{1}&\ldots&y_{n}\end{bmatrix}^{\top}$, where
$y_{1}$ $=S(x_{0}+u_{0})$
$y_{2}$ $=S(Wx_{0}+Wu_{0}+x_{1}+u_{1})$
$y_{3}$ $=S(W^{2}x_{0}+Wx_{1}+W^{2}u_{0}+Wu_{1}+x_{2}+u_{2})$
@@
$y_{T-1}$ $=S(W^{T-1}x_{0}+\cdots+x_{T-1}+W^{T-1}u_{0}+\cdots+u_{T-1}),$
where $S$ is the selector matrix for $\mathcal{A}$
From this representation we get the following for a fixed pair of neighboring streams of data.
@@
Consider the neighboring sets of data-streams $D$ and $D^{\prime}$ that change at most by one node’s contribution (let it be node $j\in[n])$, such that each for each $t\in[T]$, where $T$ denotes the total number of iterations, it holds that $x_{t}^{j}-{x^{\prime}}_{t}^{j}=c_{t}e_{j}$, where $\left|c_{t}\right|\leq 1$. Then, the node $j$ is $(\varepsilon,\delta)$-DP from the point-of-view of node $i$, where $(\varepsilon,\delta)$ is bounded by the $(\varepsilon,\delta)$-distance between the two multivariate mechanisms
@@
$\widetilde{x}_{T}=H_{T}\left(\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{T}\end{bmatrix}\otimes e_{j}\right)\quad\textrm{and}\quad H_{T}:=\begin{bmatrix}S&0&0&\dots&0\\ SW&S&0&\dots&0\\ SW^{2}&SW&S&\dots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ SW^{T-1}&SW^{T-2}&\dots&SW&S\end{bmatrix}$ (3.2)
@@
Consider the Gossip averaging algorithm (3.1) and suppose $D,D^{\prime}$ are neighboring datasets such that $D\simeq_{j}D^{\prime}$. Let $\widetilde{x}_{T}$ and $H_{T}$ be given as in Eq. (3.2). Let $H_{T}=U\Sigma V^{\top}$ be the compact SVD of $H_{T}$. Let $\widetilde{x}_{T}$ be of the form (3.2) and denote $c=\begin{bmatrix}c_{1}&c_{2}&\ldots&c_{T}\end{bmatrix}^{\top}\in\mathbb{R}^{T}$. Denote
$\Delta^{T}_{j\rightarrow i}=\max_{c\in\{-1,1\}^{T}}\left\lVert H_{T}^{+}H_{T}(c\otimes e_{j})\right\rVert_{2}=\max_{c\in\{-1,1\}^{T}}\left\lVert V^{\top}(c\otimes e_{j})\right\rVert_{2}\leq(\mathbf{1}_{T}\otimes e_{j})^{\top}\left|VV^{\top}\right|(\mathbf{1}_{T}\otimes e_{j}),$